# Do LLMs Really Think Step-by-step In Implicit Reasoning?

### v3

@@ 2.1 Expriment Design
The model will direct output the answer after our prompt, thus we take the last token of the prompt as our main research object and record its hidden states of each layers. Then, we adopt a typical linear probing method using an 1-layer MLP, to predict each of the intermediate results from the hidden states of each layer. We control the values of all intermediate results and the final answer within the range of 0-19, so that the probe is a 20-class classifier (each value corresponds to one class).
@@ 2.3 Result of Modifying the Problem Presentation
By contrast, from the results of implicit CoT in Table 1, we can clearly see that, compare to the original problems, the modified problems significantly degrade the performance for both situations (prompted or trained), especially when the order of equations is reversed. For the trained way, despite it truly thinks step by step, the accuracy even drops more significantly. The probing results in Appendix B show that reversing makes mistral-internal-CoT unable to calculate the intermediate results at all. This finding indicates implicit CoT is usually much less reliable and stable.
@@ Appendix A Difference between implicit and explicit CoT
Appendix B Additional Results
When the order of the equations in the problem is reversed, the probing results is shown in Figure 2. We can see that the intermediate results as well as the final result cannot be calculated.

### v4

@@ 2.1 Expriment Design
The model will direct output the answer after our prompt, thus we take the last token of the prompt as our main research object and record its hidden states of each layers. Then, we adopt a typical linear probing method using an 1-layer MLP, to predict each of the intermediate results from the hidden states of each layer. We control the values of all intermediate results and the final answer within the range of 0-19 (inlucding 0 and 19), so that the probe is a 20-class classifier (each value corresponds to one class). More details about why we choose this range and how we generate samples are in Appendix B.
@@ 2.3 Result of Modifying the Problem Presentation
By contrast, from the results of implicit CoT in Table 1, we can clearly see that, compare to the original problems, the modified problems significantly degrade the performance for both situations (prompted or trained), especially when the order of equations is reversed. For the trained way, despite it truly thinks step by step, the accuracy even drops more significantly. The probing results in Appendix C show that reversing makes mistral-internal-CoT unable to calculate the intermediate results at all. This finding indicates implicit CoT is usually much less reliable and stable.
@@ Appendix A Difference between implicit and explicit CoT
Appendix B Data Generation
In the probing experiment, to train a fair probe model, we need to keep the number of each label is nearly the same, i.e., they obey discrete uniform distribution.
In some our early probing experiments, we found LLMs are worse at handling negative numbers than positive numbers: the probing accuracy of the intermediate result of a negative number is generally much lower than that of a positive number. Therefore, for fairness, we make all the intermediate results $R$ random positive numbers or 0. Specifically, we make it obey this distribution:
$P(R_{i}=x)=1/(M+1),x\in[0,M],M\in N_{+}$
Here $R_{i}$ represents the result of the $i$-th step. $R_{0}$ represents the initial value. And we let $A_{i}$ represent the addend in the $i$-th step.
To generate uniform distributed $R_{i}$, in practice, we actually first generate the intermediate results using random.randint(), and then calculate $A_{i}$ by $A_{i}=R_{i}-R_{i-1}$.
We set $M$ to 19 in our experiment, because we find setting $M$ to 10 will make the data lack diversity, while a too large $M$ will slow the training of the probe.
Figure 3: The accuracy of probing the results of each step in multi-step arithmetic problems from the hidden states of mistral-internal-CoT, when the order of the equations is reversed.
Appendix C Additional Results
When the order of the equations in the problem is reversed, the probing results is shown in Figure 3. We can see that the intermediate results as well as the final result cannot be calculated, while only the initial value can be probed. This results demonstrate current implicit-CoT methods are susceptible to the format of the problem.